A university was interested in examining the overall effectiveness of its online statistics course, along with the effectiveness of particular aspects of the course. First, the university wanted to see whether the online course was better than a standard course. Second, the university wanted to know whether students learned best using Excel, using Minitab, or using no statistical package at all. The university randomly selected a group of 30 students and administered one of the different variants of the course (i.e., traditional or online, coupled with one of the software options) to each student. The success of each variant was measured by the students' average improvement between a pre-test and a post-test.

The response variable is:

  • a)the students' mean post-test score.
  • b)the difference between the students' mean pre-test and post-test scores.
  • c)an evaluation of course quality provided by an independent rate.

If we would like the treatments to be balanced in terms of the number of subjects, how many students should we assign to each treatment?

  • a)3
  • b)4
  • C)5
  • d)6 (this one is not correct)
  • e)7
    Answer #1

    1)

    Since we know that success was measured by students's average improvement between a pre-test and post-test

    = average score post test - average score pre test

    answer is b)

    2)

    Number of courses = 2

    Number of Software = 3

    Total number of treatment = 2*3 =6

    Now we have 30 students and 6 treatment which can be assigned = 30/6 = 5 students for each treatment

    answer is C)